OpenStudy (anonymous):

How do I solve this? 2x+10y=80 2x-5y=5 Help please!!

OpenStudy (amistre64):

you can subtract one from the other to get rid of worrying about one of the variables

OpenStudy (amistre64):

(2x+10y=80) -(2x - 5y = 5) --------------- 15y = 75 solve fpr y; once you know why, use either equation to solve for x